OpenStudy (anonymous):

How many grams of carbon dioxide is formed if 5000 moles of methane is burned?

OpenStudy (sk8teroy):

Equation of burning methane: \[CH_4 + 2O_2 \rightarrow CO_2 + 2H_2O\] According to the equation, if 5000 moles is burned, then 5000 moles of carbon dioxide is burned because it is at 1:1 ratio. To find the mass of carbon dioxide, we use: \[mass = mol \times M_r\] where Mr is the molecular mass. Therefore, substituting the values of the no. of moles of carbon dioxide and the molecular mass of carbon dioxide, we have: \[mass = 5000 \times 44\] \[mass of CO2 = 220000 g = 2.2 \times 10^5 g = 220 kg\]